The new calibration matrix has been tested today. The work started from 15:02, 03, Aug, 2023 UTC. Two SQZ automation was put in pause.
For the test of calibration matrix, GR_AA and BPC2 were opened. The test was done between 15:10 and 15:30.
For the measurement of actuation matrix, only BPC2 was open. Only GR_M5_X, GR_M5_Y, and GR_M7_X were tuned between 15:30 and 16:00. (Due to some unlock of FC)
Before left EGO, the system was recovered to automation exec.
The GR_M7_Y was tuned between 17:40 and 18:02.
After all these work. The filter cavity is put back to automation exec.